[figure: 9.7 cm]

MAGNETIC AND ELECTRIC TRANSPORT PROPERTIES OF Fe-Hf-O FILMS PREPARED BY SUPERSONIC PLASMA JET

METHOD

V. Soyka and L. Kraus Institute of Physics ASCR, Na Slovance 2, CZ-182 21 Praha 8, Czech Republic

Fe-Hf-O films were prepared by supersonic plasma jet deposition technique on glass and silicon substrates. The material of the nozzle is sputtered by the RF plasma generated in the Ar + H_{2} gas, incoming with the supersonic speed through the nozzle into the vacuum chamber. Composition of the films varied in the limits: 15-53 at.% of Fe, 1-14 at.% of Hf, 42-80 at.% of O. In the two-nozzle-method the nozzles are made of Fe and Hf. The film composition is controlled by the distance of Hf-nozzle from the substrate. The films show weak α-Fe peaks on X-ray diagram. The Fe_{52.75}Hf_{1.9}O_{45.35} film shows the largest value of saturation magnetisation (J_{s} = 0.56 T), determined from FMR measurements. The temperature dependence of resistivity is similar to that observed in amorphous metals. The one-nozzle-method uses a single nozzle composed of iron and hafnium rings. In this case the substrates are water-cooled. Different levels of RF signal and continuous and pulse regimes of power generator are used. The films are X-ray amorphous and their composition depends on Fe/Hf ratio in the nozzle and the preparation conditions. Samples exhibit lnR(T)∝T^{x} temperature dependence of resistance with x close to -1/2, which indicates on electron spin dependent scattering mechanism of resistivity.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : poster

Corresponding author : V. Soyka

Address for correspondence : Na Slovance2, CZ-18221 Praha 8, Czech Republic

Email address : soyka@fzu.cz